Currently Tom is 10 years older than Mike. In 5 years the sum of their ages will be 44. How old are Tom and Mike currently?
Answer by edjones(8007)

Let t=tom's current age and m=mike's
t=m+10 Now.
.
t+5+m+5=44 In five yrs.
t+m=34
m+10+m=34 Substitute m+10 for t
2m=24
m=12
t=22
.
Ed